OpenStudy (anonymous):

So confused :( The point (5/13,y) in the fourth quadrant corresponds to angle θ on the unit circle. sec θ = cot θ = help!!!!!

Nnesha (nnesha):

first number in order is represent cosine and 2nd one is sin (cosine , sin) so according to this info Cos =5/13 make a triangle in 4th quadrant and \[\rm sin = \frac{ opposite }{ hypotenuse }~~~~ \cos = \frac{ adjacent }{ hypotenuse } ~~\tan= \frac{ opposite }{ adjacent }\]

OpenStudy (anonymous):

trigonometry!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

:( i swear i have never seen this before and it is in our review packet im in alg2

Nnesha (nnesha):

so cosine = adjacent over hypotenuse use Pythagorean theorem to find 3rd side of right triangle remember triangle it's in 4th quadrant so cosine should be negative